My math teacher, unfortunately, never taught me how to do this kind of supposed "mixture" problem: With a tailwind, a jet flew 2000 miles in 8 hours, but the return trip against the wind required 10 hours. Find the jet's speed and the wind speed.
Answer by stanbon(75887)

With a tailwind, a jet flew 2000 miles in 8 hours, but the return trip against the wind required 10 hours. Find the jet's speed and the wind speed.
----------------------
With tailwind DATA:
Distance = 2000 miles ; Time = 8 hrs. ; Rate = 2000/8 = 250 mph
------------------
Return trip DATA:
Distance = 2000 miles ; Time = 10 hrs. ; Rate = 2000/10 = 200 mph
-------------------------------------
Let p be plane speed with no wind; Let w be the speed of the wind.
EQUATIONS:
p+w = 250 mph
p-w = 200 mph
---------
Add the two equations to solve for "p":
2p = 450
p = 225 mph (speed of the plane with no wind)
------------
Subsitute to solve for w:
225 + w = 250
w = 25 mph (speed of the wind)
================================
Cheers,
Stan H.